OpenStudy (anonymous):

ok I am using webassign as the math program for algebra class, and it seems whatever answer I put it it says its wrong. so try this problem for me though, its rational exponents. b^1/4 over b^5/4

OpenStudy (amistre64):

1 -------- b^(5-1/4) 1/b

OpenStudy (anonymous):

I got -1..that must be wrong

OpenStudy (amistre64):

b^(1/4) 1 1 ------- = ----------- = ------- = 1/b^1 = 1/b b^(5/4) b^[(5-1)/4] b^(4/4)

OpenStudy (amistre64):

1/b and b^(-1) are the same thing; but many programs dont want a -exponent

OpenStudy (anonymous):

Yeah I can see that now. Thank you soo much for explaining.

OpenStudy (amistre64):

youre welcome :)

OpenStudy (amistre64):

whenever you have a -exp...put it to the other side of the fraction bar and change it to +

OpenStudy (amistre64):

like this: b^-2 1 ---- = --- 1 b^2

OpenStudy (anonymous):

okay thank you I did 1/b as the answer and it liked it.